# Convolutional Monge Mapping between EEG Datasets to Support Independent Component Labeling

### v1

@@ II-B Proposed Methodology
where $\mathrm{IRFFT}_{M}$ is the inverse FFT for real-valued signals that operates on the non-negative frequencies. As discussed in Section II-A, by equalizing the channel-average PSD, this filter solves the optimal transport problem between zero-mean Gaussian distributions with circulant covariance matrices that commute and are diagonalized by the discrete-time Fourier transform. For this filter, the source PSD $\bar{\bm{p}}^{\text{S}}$ can be either the barycenter of the source dataset $\bar{\bm{p}}_{\text{S}}$ (19), its $\ell_{1}$-normalized version $\tilde{\bm{p}}_{\text{S}}$ (20), or the closest matching source signal $\bar{\bm{p}}_{*}^{\text{S}}$ (21).
II-C Application of CMMN to pre-trained IC classifiers
@@
While the proposed methodology is general and could be applied to normalize any target dataset for classifiers pre-trained on a source dataset, we focus on classifiers for automatically labeling independent components of EEG into brain and other classes [17, 7, 2]. In particular, we build on prior work [5], applying the CMMN approach to different types of classifiers (trained without CMMN):
  1. 1.
    Random forest classifiers that use features from the PSD and autocorrelation sequence of the independent components, as in the MNE-ICALabel implementation [18], a Python port of ICLabel [2].
  2. 2.
    Random forest classifiers that use the bag-of-waveforms (BoWaves) features [5], which are built on the occurrence of a dictionary of waveforms learned through a shift invariant $k$-means algorithm [19].
III Experiments and Results
III-A Data
We use two datasets for this domain adaptation study, the Imagined Emotion dataset from [8] as the source dataset and the Cue dataset from [6] as the target dataset. Both datasets include manually labeled independent components (ICs). We hereafter refer to the Imagined Emotion dataset as simply the Emotion dataset.
The Emotion dataset has data from 32 subjects (13 male and 19 female, with an age mean and standard deviation of 25.5 $\pm$ 5 years). The EEG data has 180-232 channels and sampled at 256 Hz with durations ranging from 54 to 136 minutes. Channels with abnormal activity before performing ICA were discarded, leaving 134–235 channels per subject. This dataset was collected in the US, leaving a characteristic 60 Hz spike from AC electrical wiring as seen in Fig. 1. The dataset includes 935 expert-labeled ICs in three categories: brain (570), muscle (306), and eye (59). We used 27 subjects for training and 7 for testing. The training set contains 5,786 total ICs.
Figure 1: Channel-averaged power spectral densities for each subject in the Emotion (source) dataset. Notice that some subjects are outliers in terms of overall amplitude.
@@ III-A Data
The Cue dataset has data from 12 subjects (10 male, 2 female, with an age range of 21 to 25 years), using 64 channels during data collection. This dataset’s recordings were 56 to 66 minutes long and were collected at 500 Hz, which we down-sampled to 256 Hz before performing domain adaptation. This dataset was collected in Europe, leaving a characteristic 50 Hz spike as seen in Fig. 3. Expert annotations are available for 389 ICs: brain (261), muscle (102), eye (22), and heart (4). This dataset serves as the target for cross-dataset evaluation, testing generalization of the classifiers trained on the Emotion training data.
Note that in comparison to the experiments in the original CMMN paper [1], our data contains different numbers of channels (134–235 vs. 64). Therefore, the use of channel-averaged PSDs for CMMN is essential. For further details on both datasets, please refer to [8, 6].
III-B IC classifiers
We trained multi-class classifiers using each feature set (BoWaves or PSD+Autocorrelation) on the subset of 27 subjects from Emotion dataset, with hyperparameters selected by leave-one-subject-out (LOO) cross-validation accuracy. The classifier is trained on segments of length $l_{\text{train}}$ (a hyperparameter), and validated (using the F1 score on the brain class) on segments of length $l_{\text{val}}\in\{5,50\}$ minutes, with $l_{\text{train}}\leq l_{\text{val}}$. Given the best LOO hyperparameters, a new model is trained on the entire training set of 27 subjects. Here we report test results on segments of length equal to the validation segment length, from a disjoint test set of 7 source Emotion subjects and all the target subjects in the Cue dataset.
We benchmark our method against ICLabel [2]. As a model trained on an extensive dataset of expert- and crowd-labeled ICs, ICLabel serves as a strong baseline that can be applied to both the source and target datasets without our proposed CMMN. Importantly, ICLabel uses the spatial information in the form of the mixing vector corresponding to the topographic map for each IC, which is not included in the BoWaves or the PSD and Autocorrelation feature sets.
The source code for our implementation and experiments is publicly available at https://github.com/cniel-ud/ICWaves.
III-C Results
@@
We show that the normalizing filters learned display intuitive mapping behaviors. The Emotion dataset [8] was collected in the US, and displays a characteristic 60 Hz line noise as seen in Fig. 1. The Cue dataset [6] was collected in Europe, and displays a characteristic 50 Hz line noise as seen in Fig. 3; this is an artifact of the two regions’ different electrical regimes. When mapping from the Europe dataset to the US dataset, we can see in Fig. 4 that the filters learned to minimize the 50 Hz spike and introduce a spike at 60 Hz.
Figure 2: This is the $\ell_{1}$-normalized barycenter computed from the source Emotion dataset. In the Barycenter mapping scheme, all target signals are filtered such that their robust channel-average PSD matches this. Notice the large spike at 60 Hz.
Figure 3: Here we show the PSD for different subjects in the target Cue dataset. Notice the spike at 50 hz.
@@
Figure 4: Here we see the frequency response of filters learned for mapping the Cue dataset to the Emotion dataset using the Barycenter mapping scheme with a normalized barycenter. Notice that the noise at 50 Hz is reduced and the noise at 60 Hz is relatively amplified. Overall, the filter attenuates since the normalized barycenter has lower magnitudes.
Figure 5: Similarly to Fig. 4, we see the filters learned for mapping the Cue dataset to the Emotion dataset, but here they are learned using the Subj-to-subj mapping scheme. Even though the mapping scheme is different, the line noises are visibly still being ‘swapped’.
TABLE I: F1 score of brain-labeled independent components for different test segment lengths and CMMN schemes. Best performance for each segment length is bolded, and best scheme is underlined. The last column contains the p-value for a one-sided Wilcoxon rank-sign test between no filtering and CMMN filtered (subject-to-subject for BoWaves, and normalized barycenter for PSD/Autocorr.) across the 12 subjects.
Classifier Length (minutes) No filtering Barycenter $\ell_{1}$-Norm. barycenter Subj-to-subj p-value
BoWaves 5 0.88 $\pm$ 0.05 $8±\pm0$.04 0.87 $\pm$ 0.05 0.90 $\pm$ 0.08 0.0593
PSD/Autocorr. 5 0.74 $\pm$ 0.10 -b 0.35 $\pm$ 0.30 0.81 $\pm$ 0.10 -b
[…]
ICLabel 5 0.88 $\pm$ 0.06 N/Ac N/A N/A N/A
BoWaves 50 0.89 $\pm$ $90±\pm 0.0$8 0.88 $\pm$ 0.07 0.91 $\pm$ 0.08 0.0412
PSD/Autocorr. 50 0.79 $\pm$ 0.10 -b 0.40 $\pm$ 0.28 0.84 $\pm$ 0.12 -b
PSDa/Autocorr. 50 0.83 $\pm$ $\pm0.09 0.86$ $\pm$ 0.08 0.85 $\pm$ 0.17 0.1696
ICLabel 50 0.89 $\pm$ 0.05 N/A N/A N/A N/A
aUsing 0-1 range normalization rather than max absolute value normalization.
bAfter noticing the issues with the max absolute value normalization we did not run this case.
cWe did not apply ICLabel to the CMMN normalized signals, as it was trained over a wide variety of datasets and should be already invariant.
As seen in Table I, our results show an improvement compared to no filtering when using Emotion-trained classifiers on the Cue dataset for labeling ICs using an appropriate CMMN scheme for the classifier. Also unfiltered performance of the BoWaves classifier matches that of ICLabel, indicating that the spatial information in the topographic map feature is not necessary.
The subject-to-subject mapping scheme improves all of the classifier’s performance. For both segment lengths, the subject-to-subject mapping scheme improves the BoWaves performance above the ICLabel benchmark. For classifiers that use the PSD and autocorrelation features, the subject-to-subject mapping scheme improves the brain class F1-score from 0.74 to 0.81 at 5 minutes (+0.07) and from 0.79 to 0.84 at 50 minutes (+0.05).
The performance of the PSD/Autocorrelation classifier was highly dependent on the feature normalization method, particularly for the barycenter mapping schemes. Our initial approach, taken from ICLabel [2], normalized the decibel-scaled PSD by its maximum absolute value. This method proved unsuitable for the $\ell_{1}$-normalized barycenter CMMN schemes, which uses $\ell_{1}$-normalized PSDs and thus have a much lower overall signal scale compared to the single-subject PSDs used in the subject-to-subject scheme. To resolve this sensitivity to signal scale, we switched to a more robust min-max normalization that scales features to a 0–1 range (denoted as PSDa/Autocorr. in Table I). This change led to a significant performance increase for the barycenter methods. For instance, with the $\ell_{1}$-normalized barycenter, the F1-score for 5-minute segments improved from 0.77 to 0.84, demonstrating the effectiveness of CMMN with the corrected feature scaling.
We perform a hypothesis test whether the filtering (subject-to-subject for BoWaves, and normalized barycenter for PSD/Autocorrelation) has a significant improvement in F1 score for the brain-labeled ICs. We use a Wilcoxon rank sign test ($n=12$, one-sided alternative), and find significant improvements at a level of 0.05 for two of the four cases marked in the table.
[…]
@@
TABLE II: F1 score of brain-labeled independent components of classifiers within domain on held-out subjects.
@@
IV RELATED WORKS
Domain adaptation is a well-studied problem, with optimal transport methods providing principled approaches [20]. In the context of EEG, recent work [21] uses a geodesic optimization approach to tackle domain adaptation in both the data and label space. Other work [22, 23] use Maximum Mean Discrepancy (MMD) [24] to align source and target EEG data distributions. While MMD is a powerful and general purpose divergence, the Bures-Wasserstein distance that underlies CMMN is sufficient for zero-mean stationary Gaussian processes.
A clustering-based approach was previously applied [25] to help tackle high variability between subjects. We note that the subject-to-subject and barycenter approaches are two extreme clusterings, and an intermediate clustering could be investigated.
Denoising EEG through independent component analysis is well-known [26, 27] and the automatic labeling of ICs with machine learning has seen a number of advancements [17, 7, 2, 5, 28]. An early method [17] used only six features to reach high levels of inter-expert agreement on labeling of independent components. We note that prior work [7] also classified independent components from the same two datasets [6, 8] that we use, using 65 features. In comparison, BoWaves and PSD/Autocorr use larger number of features characterizing the time-series in terms of its temporal and spectral characteristics. Finally, ICLabel [2] has since become a standard for automatic labeling of independent components; thus, we use it as a benchmark of our classifiers. It is noteworthy that with the proposed CMMN the BoWaves classifier outperforms ICLabel without any access to the spatial information.
V Discussion
One possible concern is whether channel-averaging is a sufficient characterization of multi-channel EEG given the spectral differences in signals across the scalp. Of course, if the location of the electrodes on the scalp is vastly different such that the spectral content is not comparable, then channel-averaged PSD may not be meaningful, but for EEG montages with whole scalp this should not be an issue.
An interesting extension is to consider domain adaptation for target data collected from EEG montages with limited scalp coverage. In these cases, the channel-averaging in the source should be performed on a subset of channels with the same scalp coverage. Again, as the classifiers tested here only depend on spectral or temporal features, the independent component classifiers will not be limited by the channel coverage. This is because the brain ICs seen during training likely occurred across the brain.
VI Conclusion
In this work we made several key extensions to the recently introduced CMMN methodology [1] to enable domain adaptation of BoWaves classifiers [19, 5, 28] between EEG datasets for independent component classification for artifact removal. We introduce filters defined by channel-averaged PSDs along with a subject-to-subject mapping scheme, and show that our CMMN method results in improvements when classifying brain versus non-brain independent components, achieving domain adaptation between two different datasets with significant differences. With the domain adaptation the BoWaves classifiers exceed the performance of the popular ICLabel [2] classifier. Our method advances the work on EEG artifact-removal across distinct datasets, an area that is crucial for increasing the clinical utility of EEG recordings.
ACKNOWLEDGMENT

### v2

@@ 1 Methodology
where $\mathrm{IRFFT}_{M}$ is the inverse FFT for real-valued signals that operates on the non-negative frequencies. As discussed in Appendix A, by equalizing the channel-average PSD, this filter solves the optimal transport problem between zero-mean Gaussian distributions with circulant covariance matrices that commute and are diagonalized by the discrete-time Fourier transform. The only difference between the barycenter and subject-to-subject formulations is the choice of Gaussian distribution for the source. That is, the source PSD $\bar{\bm{p}}^{\text{S}}$ can be either the barycenter of the source dataset $\bar{\bm{p}}_{\text{S}}$ as in (1), the $\ell_{1}$-normalized barycenter $\tilde{\bm{p}}_{\text{S}}$ as in (1), or the closest matching source signal $\bar{\bm{p}}_{*}^{\text{S}}$ as in (2).
1.1 Application of CMMN to pre-trained IC classifiers
@@
While the proposed methodology is general and could be applied to normalize any target dataset for classifiers pre-trained (without CMMN) on a source dataset, we focus on classifiers for automatically labeling independent components of EEG into brain and other classes [18, 5, 14]. In particular, we build on recent work for labeling ICs without spatial information [11], applying the CMMN approach to random forest classifiers pre-trained on two different feature vectors: the PSD  [^2] In contrast to the original ICLabel implementation, we use a min-max normalization of the log-scale PSD instead of dividing by its max-absolute value because it is more robust to the scale of the EEG signals and provided a better classification performance. and autocorrelation sequence of the independent components, as in the MNE-ICALabel implementation [10], a Python port of ICLabel [14].
2 Experiments and Results
@@
We use two datasets for this domain adaptation study, the Imagined Emotion (Emotion for short) dataset from [12] as the source dataset and the Cue dataset from [8] as the target dataset. Both datasets include manually labeled independent components (ICs). The Emotion dataset has data from 32 subjects (13 male and 19 female, with an age mean and standard deviation of 25.5 $\pm$ 5 years). The EEG data has 180-232 channels and sampled at 256 Hz with durations ranging from 54 to 136 minutes. The dataset includes 935 expert-labeled ICs in three categories: brain (570), muscle (306), and eye (59). We used 27 subjects for training and 7 for testing. The training set contains 5,786 total ICs. The Cue dataset has data from 12 subjects (10 male, 2 female, with an age range of 21 to 25 years), using 64 channels during data collection. This dataset’s recordings were 56 to 66 minutes long and were collected at 500 Hz, which we down-sampled to 256 Hz before performing domain adaptation. Expert annotations are available for 389 ICs: brain (261), muscle (102), eye (22), and heart (4). This dataset serves as the target for cross-dataset evaluation, testing generalization of the classifiers trained on the Emotion training data. Note that in comparison to the experiments in the original CMMN paper [7], our data contains different numbers of channels (134–235 vs. 64). Therefore, the use of channel-averaged PSDs for CMMN is essential. For further details on both datasets, please refer to [12, 8].
@@
We show that the normalizing filters learned display intuitive mapping behaviors. The Emotion dataset [12] was collected in the US, and displays a characteristic 60 Hz line noise as seen in Fig. 2 in the appendix. The Cue dataset [8] was collected in Europe, and displays a characteristic 50 Hz line noise as seen in Fig. 3 in the appendix. When mapping from the Europe dataset to the US dataset. We can see in Fig. 1 that the filters learned to minimize the 50 Hz spike and introduce a spike at 60 Hz.
@@
Figure 1: Frequency response of filters learned for mapping the Cue dataset to the Emotion dataset. (Left) Barycenter mapping scheme with a $\ell_{1}$-normalized barycenter. Notice that the noise at 50 Hz is reduced and the noise at 60 Hz is relatively amplified. Overall, the filter attenuates since the normalized barycenter has lower magnitudes.(Right) Subj-to-subj mapping scheme. Even though the mapping scheme is different, the line noises are visibly still being ‘swapped’.
We trained multi-class classifiers using both PSD and Autocorrelation features on the subset of 27 subjects from Emotion dataset, with hyperparameters selected by leave-one-subject-out (LOO) cross-validation. The classifier is trained on segments of length $l_{\text{train}}$ (a hyperparameter), and validated (using the F1 score on the brain class) on segments of length $l_{\text{val}}\in\{5,50\}$ minutes, with $l_{\text{train}}\leq l_{\text{val}}$. Given the best LOO hyperparameters, a new model is trained on the entire training set of 27 subjects. Here we report test results on segments of length equal to the validation segment length, from a disjoint test set of 7 source Emotion subjects and all the target subjects in the Cue dataset. We benchmark our method against ICLabel [14]. As a model trained on an extensive dataset of expert- and crowd-labeled ICs, ICLabel serves as a strong baseline that can be applied to both the source and target datasets without our proposed CMMN. We did not apply ICLabel to the CMMN normalized signals, as it was trained over a wide variety of datasets and should be already invariant. The source code for our implementation and experiments is publicly available at https://github.com/cniel-ud/ICWaves.
The performance on held-out subjects of the source Emotion dataset is shown in Table 1. PSD/Autocorr has the best performance at both time lengths.
@@
Table 1: F1 score of brain-labeled independent components of classifiers within domain on held-out subjects.
@@
Table 2: F1 score of brain-labeled independent components for different test segment lengths and CMMN schemes. Best performance for a classifier at a segment length is underlined and best classifier for a segment length is bolded. The last column contains the p-value for a one-sided Wilcoxon rank-sign test between no filtering and CMMN filtered (normalized barycenter for PSD/Autocorr.).
Classifier Len. No filtering Barycenter $\ell_{1}$-norm. Bary. Subj-to-subj p-value
@@
ICLabel 5 0.88 $\pm$ 0.06
PSD/Autocorr. 50 0.83 $\pm$ $\pm0.09 0.86$ $\pm$ 0.08 0.85 $\pm$ 0.17 0.1696
ICLabel 50 0.89 $\pm$ 0.05
As seen in Table 2, our results show that, when compared to no filtering, using the appropriate CMMN scheme improves the performance of Emotion-trained classifiers on the Cue dataset. The subject-to-subject mapping scheme improves the brain class F1-score from 0.77 to 0.79 at 5 minutes and from 0.83 to 0.85 at 50 minutes. With the $\ell_{1}$-normalized barycenter, the PSD/Autocorr classifiers’ F1-score improved from 0.77 to 0.84 at 5 minutes, and from 0.83 to 0.86 at 50 minutes. We use a Wilcoxon rank sign test ($n=12$, one-sided alternative) to test whether the filtering (normalized barycenter) has a significant improvement in F1 score for the brain-labeled ICs, and find significant improvement at a level of 0.05 for the 5 minute length.
3 Conclusion
In this work we made key extensions to the recently introduced CMMN methodology [7] to enable domain adaptation between EEG datasets for independent component classification for artifact removal. We introduce filters defined by channel-averaged PSDs along with a subject-to-subject mapping scheme, and show that our CMMN method results in improvements when classifying brain versus non-brain independent components, achieving domain adaptation between two different datasets with significant differences. Our method advances the work on EEG artifact-removal across distinct datasets, an area that is crucial for increasing the clinical utility of EEG recordings.
Acknowledgments and Disclosure of Funding
@@ 3 Conclusion
Appendix A Background on Convolutional Monge Mapping Normalization
The Wasserstein-2 distance between Gaussian distributions has a closed form [3, 6] known as the Bures-Wasserstein distance or Fréchet distance, expressed in terms of the means $\bm{m}^{\text{S}},\bm{m}^{\text{T}}$ and covariance matrices $\bm{\Sigma}^{\text{S}},\bm{\Sigma}^{\text{T}}$, as
where the second term in the square root is the squared Bures distance. Under the stationarity assumption, a discrete-time zero-mean Gaussian process is completely described by its auto-covariance matrix, which is a symmetric and Toeplitz matrix formed from the auto-correlation sequence $r[k]=\mathbb{E}[x[n]x[n+k]]$. Assuming a sufficiently long truncation of $r[\tau]$ that yields a $N\times N$ circulant matrix that is positive-semidefinite, then the discrete Fourier transform (DFT) matrix $\bm{F}$ diagonalizes it, $\bm{F}\bm{\Sigma}\bm{F}^{H}=\mathrm{diag}(\bm{\lambda})$, where $\bm{F}^{H}$ denotes the Hermitian transpose of $\bm{F}$. The resulting spectrum $\bm{\lambda}=\mathrm{diag}(\bm{F}\bm{\Sigma}\bm{F}^{H})$ corresponds to an statistical estimate of the power spectral density (PSD). The squared Wasserstein-2 distance between two discrete-time, zero-mean stationary Gaussian random processes with auto-covariance $\bm{\Sigma}^{\text{S}}$ and $\bm{\Sigma}^{\text{T}}$, which are circulant, simplifies to
$W_{2}(\mathcal{N}(\bm{0},\bm{\Sigma}^{\text{S}}),\mathcal{N}(\bm{0},\bm{\Sigma}^{\text{T}}))=\lVert\bm{\lambda}^{\text{S}\odot\frac{1}{2}}-\bm{\lambda}^{\text{T}\odot\frac{1}{2}}\rVert_{2}=\sqrt{\sum_{n=0}^{N-1}\left(\sqrt{\lambda^{\text{S}}[n]}-\sqrt{\lambda^{\text{T}}[n]}\right)^{2}}.$ (6)
Assuming all covariance matrices are circulant and positive semidefinite, it is straightforward to show that $\bm{\Sigma}_{\text{S}}=\bm{F}^{H}\mathrm{diag}(\bar{\bm{\lambda}}_{\text{S}})\bm{F}$, where $\bar{\bm{\lambda}}_{\text{S}}=(\frac{1}{I}\sum_{i}\bm{\lambda}_{i}^{\text{S}\odot\frac{1}{2}})^{\odot 2}$.
$\bm{\lambda}^{y}$ $=\mathbb{E}[|\bm{F}\bm{y}|^{\odot 2}]=\mathbb{E}[|(\bm{F}\bm{x})\odot\bm{H}|^{\odot 2}]=\mathbb{E}[|\bm{F}\bm{x}|^{\odot 2}\odot|\bm{H}|^{\odot 2}]=|\bm{H}|^{\odot 2}\odot\bm{\lambda}^{\text{T}},$ (8)
Clearly, the distance is zero for any filter $h[n]$ such that $|\bm{H}|=\bm{\lambda}^{\text{S}\odot\frac{1}{2}}\odot\bm{\lambda}^{\text{T}\odot-\frac{1}{2}}\in\mathbb{R}_{\geq 0}^{N}$, i.e., where the magnitude of the frequency response is the square-root of the power spectral densities $|H[n]|=\sqrt{\frac{{\lambda}^{\text{S}}[n]}{\lambda^{\text{T}}[n]}},n\in\{0,\ldots,N-1\}$.
Appendix B Additional Results
Figure 2: (Top) Channel-averaged power spectral densities for each subject in the Emotion dataset. Notice that some subjects are outliers in terms of overall amplitude. (Bottom) This is the $\ell_{1}$-normalized barycenter computed from the source Emotion dataset. In the Barycenter mapping scheme, all target signals are filtered such that their robust channel-average PSD matches this. Notice the large spike at 60 Hz.
Figure 3: PSDs for different subjects in the target Cue dataset. Notice the spike at 50 Hz.